# A Lightweight Approach for State Machine Replication

### v1

@@ 2 Consensus with a Median Rule
We start with a very simple rule for stabilizing consensus that combines the median rule in [19] with undecided-state dynamics (e.g., [4]). The specific algorithm of the $(k,l)$-median rule can be found in Algorithm 1. Roughly, each server requests $k$ values from random peers, and if it receives at least $\ell$ it adopts the median value as its own, else it adopts $\bot$ (corresponding to the undecided state). Since blocked servers will never receive any replies, their value will be equal to $\bot$ at the end of the round. The validity property (Definition 1.2) is trivially satisfied by the $(k,\ell)$-median rule. Thus, it remains to show agreement and availability. Throughout the paper, we will focus on the specific case that $k=6$ and $\ell=3$ since that turns out to be the most suitable choice of parameters for our goals.
@@ 2.1 Availability
We call a server useful in round $t$ if it stores a value at the beginning of $t$ and is non-blocked during round $t$. A useless server (as in, not useful) essentially corresponds to the undecided state. By definition, availability only holds as long as a constant fraction of servers is useful. We will show three results that provide important insights into availability when using the $(6,3)$-median rule. Now we try to give some intuitive understanding of the claims and the way we will use them later on.
@@
If the initial fraction of useful servers is at least $1/2$ then for any 1-late $1/10$-blocking adversary, the fraction of useful servers monotonically converges to at least $3/4$ in just $O(\log n)$ rounds, and once it is at least $3/4$, it will be at least $3/4$ for polynomially many rounds, w.h.p.
@@
If at least a $3/10$-fraction of the servers is permanently blocked, then irrespective of the initial number of useful servers, within $O(\log n)$ rounds no server will be useful any more, w.h.p.
@@ 2.2 Agreement
To prove agreement, we show that, perhaps interestingly, applying the $(6,3)$-median rule to the set of all servers can be reduced to applying the $(3,3)$-median rule only to the set of useful servers, which allows us to disregard useless severs and brings simplifications for the subsequent analysis of the agreement property. Thus, it suffices to focus on the subsets of useful servers and to prove the following statement inspired by the analysis of the median rule in [19]. More details and a generalization to a much wider class of $(k,\ell)$-rules”, as we will show next. The $(k,\ell,f)$-rule is defined in algorithm 2.
Preconditions:
Consider any function $f:K^{\ell}\rightarrow K$ that is invariant to a reordering of its arguments and that satisfies validity, i.e., for all $x_{1},\ldots,x_{\ell}\in K$, $f(x_{1},\ldots,x_{\ell})\in\{x_{1},\ldots,x_{\ell}\}$.
@@
Selection Let $i$ be any of the non-blocked servers in round $t$ and $x^{\prime}_{i}$ be its key at the beginning of round $t+1$. For any key $x\in K$ and multiset of keys $X$ of size $|U_{t}|$ it holds:
@@
For any sequence $(U_{1},U_{2},U_{3},\dots)$ of useful servers with $n/4\leq|U_{t}|\leq n$ for all $t$, the $(3,3)$-median rule applied by $U_{t+1}$ on $U_{t}$ for all $t\geq 1$ achieves agreement in $O(\log n)$ rounds, w.h.p.
@@
For the proof of Lemma 2.6, we need the following notation. We say that a random variable $Y$ stochastically dominates a random variable $Z$, and write $Y\succeq Z$, if $Pr[Y\geq x]\geq Pr[Z\geq x]$ for any $x$. We start with a claim that follows from the proof of Claim 2.6 in [19].
@@
where $\varepsilon\to 0$ as $n_{t+1}\to\infty$. For $x\geq 0$, the value of $\Phi(x)$ can be bounded as follows (see [19]):
@@
Thus, there is an increasingly stronger drift to $\sqrt{(c\ln n)/n}$ as $\delta_{t}$ gets larger so that similarly to Lemma 2.8 in [19], the following result can be shown:
@@
As can be easily checked, this is maximized for $i=\lceil n_{t}/2\rceil$, with $g(i)$ being roughly $\frac{3}{2\cdot n_{t}}$ for such an $i$. For each value $x$ and round $t$, let us define the set of heavy identifiers $\mathcal{H}_{x,t}$ as the subset of the $\Phi_{t}=2\sqrt{cn_{t}\ln n_{t}}$ identifiers $id(x_{i})$ with $x_{i}=x$ of largest gravity (or all such values if there are less than $\Phi_{t}$), where $c>0$ is a sufficiently large constant. By definition, $0\leq|\mathcal{H}_{x,t}|\leq\Phi_{t}$. Following along the same line of arguments as in [19], the following lemma can be shown:
@@ 3.1 Broadcasting with the $(k,\ell)$-Gossip Rule
If less than $c\log n$ useful servers have $x$, then we model the number of servers that have $x$ as a random walk $X_{t}\in\{0,\ldots,q\}$ (with $q=\lceil c\log n\rceil$), where both the states 0 and $q$ are absorbing. We know from the proof of Lemma 3.1 that this random walk becomes more biased towards $q$ as we move closer to $q$, i.e., we have the probability $\Pr(X_{t+1}\geq\tfrac{3}{2}X_{t})\geq 1-e^{-c^{\prime}X_{t}}$ for some $c^{\prime}>0$. This lets us apply a claim from [19] reproduced in Appendix A Lemma A.5: after $\smash{O}(\log n)$ rounds the random walk will be in state 0 or $q$.
@@ 6 Recovery
Therefore, we require servers to maintain persistent information on their states and logs, while at the same time keeping the protocol compact, i.e., allowing servers to eventually commit commands to their shared state and subsequently forget about them. The primary challenge regarding such a mechanism is that during a surge the adversary could allow some servers to progress and commit commands while preventing others from learning these. Subsequently, if the adversary reverses this situation and permits the previously blocked servers to progress, those servers might commit a different set of commands. Such a situation is in violation of the recovery problem (see Definition 1.5). Our goal in this section is to ensure that the sequence of committed commands does not diverge so that we can quickly return to a valid solution for the commitment problem (Definition 1.4), once the adversary becomes benign again. Formally, we show the following theorem.
@@ 6.1 The Recovery Protocol
We begin with a brief overview of the recovery protocol, followed by pseudo-code (Algorithm 6). In the recovery protocol, time is divided into periods of $T$ rounds, referred to as $T$-windows. Whenever a $T$-window ends and the protocol behaves as expected, servers commit commands and create a new checkpoint that reflects the current state of servers. We define a checkpoint $C_{i}$ of server $i$ as a triple $(S,P,W)$, where $S$ is the state of the state machine, $P$ is a sequence of sufficiently old (but not yet committed) commands, and $W$ is the $T$-window index when the checkpoint was taken. The parameters $S$ and $P$ represent a valid state of server $i$ at some point, while $W$ provides a priority ordering (where newer checkpoints take precedence). We also maintain a reset variable $R_{i}$ for each server $i$, indicating whether $i$ intends to revert to a previous checkpoint. This variable assumes values reset, no-reset, or $\bot$. If $R_{i}=\texttt{no-reset}$, server $i$ expects normal operation and a benign adversary. If $R_{i}=\bot$, server $i$ believes a surge is currently in progress. Finally, if $R_{i}=\texttt{reset}$, server $i$ considers the surge over and that a rollback to a prior checkpoint is needed.
@@
The recovery protocol can be viewed as two separate protocols: the reset protocol and the checkpoint protocol that run alongside the extended median rule, and incorporate the functionality of the compact median rule with certificates. The reset protocol is governed by each server’s reset state, $R_{i}$. At the end of each $T$-window, if server $i$ has an empty log $L_{i}$ (from the extended median protocol), it concludes that it was blocked and sets $R_{i}=\texttt{reset}$. Otherwise, $R_{i}=\texttt{no-reset}$. Within a $T$-window, in every round, each server $i$ queries the reset states $R_{j}$ of $k$ randomly chosen servers. If at least $\ell$ responses are received and any of them is no-reset, the server sets $R_{i}=\texttt{no-reset}$; if all responses are reset, it sets $R_{i}=\texttt{reset}$. If fewer than $\ell$ responses are received, $R_{i}$ becomes $\bot$, and a server in the $\bot$ state does not answer queries about its own state $R_{i}$. This design ensures that if the adversary surges (i.e., blocks enough servers) during a $T$-window, the system falls into a “death spiral” where all servers eventually switch to $R_{i}=\bot$. Otherwise, the servers converge on either reset or no-reset. Finally, if a server ends a $T$-window with $R_{i}=\texttt{reset}$, it rolls back to a previously agreed-upon checkpoint.
@@
The checkpoint protocol revolves around the triple $C_{i}$. Servers exchange checkpoints by issuing $k$ random queries, analogous to how they exchange logs $\textit{L}_{i}$ or states $R_{i}$. The key distinction is that checkpoints are prioritized by age, given by the third parameter $W$. If the newest received checkpoint $(S^{\prime},P^{\prime},W^{\prime})$ has a higher number $W^{\prime}$ than its own checkpoint, it immediately adopts it and updates its state $S_{i}:=S^{\prime}$ and its log $L_{i}:=P^{\prime}$, since the server is outdated. This procedure spreads the newest checkpoints (those with the largest $W$) within one $T$-window, assuming $T$ is sufficiently large.
@@
For conciseness, we will assume in this section that the functionality of the compact median protocol with certificates is subsumed in the state machine, in particular, the required data structure centered around client sequence numbers (Section 4) and the Merkle hash forests (Section 5). The techniques shown in sections 4 and 5 can be adapted to work in the recovery protocol. Consequently, we will assume here that client commands are unique.
@@
    • $W$: number of $T$-window (initially 0, assuming the first $T$-window is 0)
@@ 7 Conclusion
We introduced a lightweight solution for state machine replication under an adversary capable of blocking servers. As shown in [19], the median rule can tolerate $O(\sqrt{n})$ adversarial servers, suggesting that our approach can be extended to handle not just blocked but also $O(\sqrt{n})$ malicious servers. Furthermore, we anticipate that it can accommodate server churn by allowing the adversary to have a $\beta$-fraction of blocked, joining or leaving servers in each round, thus removing the requirement of a static server set.
Our $1/10$ bound for blocked servers to ensure liveness is not fundamental. By tuning parameters, the protocol tolerates any 1-late $(1-\varepsilon)$-blocking adversary for constant $\varepsilon>0$. Concerning partition tolerance: if the network splits into two similarly sized groups that keep committing commands, safety would be broken. With our chosen parameters, safety even holds under any partition as all but at most one component will “spiral to death”, and liveness still holds for commands injected in the largest component provided it contains at least a $9/10$ fraction of servers. This provides a trade-off in the sense of the CAP theorem between safety, liveness, and partition tolerance [23].
Moreover, we believe our protocols can be extended so that commitment errors (for example, some command (i) is not known to all useful servers or (ii) does not have the same position at every useful server when committed, or (iii) different commands are committed for the same sequence number of some client) are not persistent. First of all, we ensure that commitment errors are very unlikely and, therefore, rare in the first place. Second, the proof of Theorem 3 implies that the extended median rule achieves agreement on the position of a command $x$ within $O(\log n)$ rounds, w.h.p., irrespective of how the logs look like initially, i.e., commitment errors only have limited effects on other commands. Third, our recovery protocol only takes another checkpoint of a shared state every $T=\Theta(\log n)$ rounds so that there is sufficient time to arrive at a consensus checkpoint that resolves disagreement on the shared state caused by a commitment error, by making use of the median rule. This pushes the probability of a persistent error to a point where it can be truly ignored.
@@ Appendix A Probabilistic Concepts
In this section, we are going to cover the probabilistic concepts and inequalities that we use frequently throughout this paper. The following two inequalities will turn out to be useful (see [32]).
@@
Finally, we state a technical lemma from from [19] (Claim 3.4), needed to show that in a $(k,\ell)$-gossip-rule, if only a single server initially has a broadcast value, then within $O(\log n)$ rounds, it will either spread to at least $c\log n$ servers or no server holds that value anymore.

### v2

@@ 2 Consensus with a Median Rule
We start with an analysis of our proposed mechanism for stabilizing consensus that combines the median rule in [DoerrGMSS11] with undecided-state dynamics (e.g., [AngluinAE08]). The specific algorithm of the $(k,l)$-median rule can be found in Algorithm 1. Roughly, each server requests $k$ values from random peers, and if it receives at least $\ell$ it adopts the median value as its own, else it adopts $\bot$ (corresponding to the undecided state). Since blocked servers will never receive any replies, their value will be equal to $\bot$ at the end of the round. The validity property (Definition 1.2) is trivially satisfied by the $(k,\ell)$-median rule. Thus, it remains to show agreement and availability. Throughout the paper, we will focus on the specific case that $k=6$ and $\ell=3$ since that turns out to be the most suitable choice of parameters for our goals.
@@ 2.1 Availability
We call a server useful in round $t$ if it stores a value at the beginning of $t$ and is non-blocked during round $t$, and otherwise it is called useless. By definition, availability only holds as long as a constant fraction of servers is useful. We will show three results that provide important insights into availability when using the $(6,3)$-median rule. The proofs are relatively technical. Here, we give an intuitive understanding of the claims and the way we use them later on.
@@
If the initial fraction of useful servers is at least $1/2$ then for any 1-late $1/10$-blocking adversary, the fraction of useful servers monotonically converges to at least $3/4$ in just $O(\log n)$ rounds, w.h.p., and once it is at least $3/4$, it will be at least $3/4$, w.h.p., for any number of rounds that is polynomial in $n$.
@@
If at least a $\tfrac{3}{10}$-fraction of servers is permanently blocked, then irrespective of the number of useful servers, within $O(\log n)$ rounds no server is useful anymore, w.h.p.
@@ 2.2 Agreement
We show that, perhaps interestingly, applying the $(6,3)$-median rule to the set of all servers can be reduced to applying the $(3,3)$-median rule only to the set of useful servers, which allows us to disregard useless severs and brings simplifications for the subsequent analysis of the agreement property. Thus, it suffices to focus on the subsets of useful servers and to prove the following statement inspired by the analysis of the median rule in [DoerrGMSS11]. More details and a generalization to a much wider class of $(k,\ell)$-rules”, as we will show next. The $(k,\ell,f)$-rule is defined in algorithm 2.
Preconditions: Consider any function $f:K^{\ell}\rightarrow K$ that is invariant to a reordering of its arguments and that satisfies validity, i.e., for all $x_{1},\ldots,x_{\ell}\in K$, $f(x_{1},\ldots,x_{\ell})\in\{x_{1},\ldots,x_{\ell}\}$.
@@
Let $i$ be any of the non-blocked servers in round $t$ and $x^{\prime}_{i}$ be its key at the beginning of round $t+1$. For any key $x\in K$ and multiset of keys $X$ of size $|U_{t}|$ it holds:
@@
For any sequence $(U_{1},U_{2},U_{3},\dots)$ of sets of useful servers satisfying $n/4\leq|U_{t}|\leq n$ for all $t$, the $(3,3)$-median rule applied by $U_{t+1}$ on $U_{t}$ for all $t\geq 1$ achieves agreement in $O(\log n)$ rounds, w.h.p.
@@
For the proof of Lemma 2.6, we need the following notation. We say that a random variable $Y$ stochastically dominates a random variable $Z$, and write $Y\succeq Z$, if $Pr[Y\geq x]\geq Pr[Z\geq x]$ for any $x$. We start with a claim that follows from the proof of Claim 2.6 in [DoerrGMSS11].
@@
where $\varepsilon\to 0$ as $n_{t+1}\to\infty$. For $x\geq 0$, the value of $\Phi(x)$ can be bounded as follows (see [DoerrGMSS11]):
@@
Thus, there is an increasingly stronger drift to $\sqrt{(c\ln n)/n}$ as $\delta_{t}$ gets larger so that similarly to Lemma 2.8 in [DoerrGMSS11], the following result can be shown:
@@
As can be easily checked, this is maximized for $i=\lceil n_{t}/2\rceil$, with $g(i)$ being roughly $\frac{3}{2\cdot n_{t}}$ for such an $i$. For each value $x$ and round $t$, let us define the set of heavy identifiers $\mathcal{H}_{x,t}$ as the subset of the $\Phi_{t}=2\sqrt{cn_{t}\ln n_{t}}$ identifiers $id(x_{i})$ with $x_{i}=x$ of largest gravity (or all such values if there are less than $\Phi_{t}$), where $c>0$ is a sufficiently large constant. By definition, $0\leq|\mathcal{H}_{x,t}|\leq\Phi_{t}$. Following along the same line of arguments as in [DoerrGMSS11], the following lemma can be shown:
@@ 3.1 Broadcasting with the $(k,\ell)$-Gossip Rule
If less than $c\log n$ useful servers have $x$, then we model the number of servers that have $x$ as a random walk $X_{t}\in\{0,\ldots,q\}$ (with $q=\lceil c\log n\rceil$), where both the states 0 and $q$ are absorbing. We know from the proof of Lemma 3.1 that this random walk becomes more biased towards $q$ as we move closer to $q$, i.e., we have the probability $\Pr(X_{t+1}\geq\tfrac{3}{2}X_{t})\geq 1-e^{-c^{\prime}X_{t}}$ for some $c^{\prime}>0$. This lets us apply a claim from [DoerrGMSS11] reproduced in Appendix A Lemma A.5: after $\smash{O}(\log n)$ rounds the random walk will be in state 0 or $q$.
@@ 6 Recovery
Therefore, we require servers to maintain persistent information on their states and logs, while at the same time keeping the protocol compact, i.e., allowing servers to eventually commit commands to their shared state and subsequently delete them. The primary challenge is that during a surge the adversary could allow some servers to progress and commit commands while preventing others from learning these. If the adversary then reverses this situation and permits the previously blocked servers to progress, those servers might commit a different set of commands, in violation of Definition 1.5. Our goal in this section is to ensure that the sequence of committed commands does not fork even under a surge and quickly return to a valid solution for the commitment problem (Definition 1.4), once the adversary becomes benign again. Formally, we show the following theorem.
@@ 6.1 The Recovery Protocol
We begin with a brief overview of the recovery protocol, followed by pseudo-code (Algorithm 6). In the recovery protocol, time is divided into periods of $T$ rounds, referred to as $T$-windows. Whenever a $T$-window ends and the protocol behaves as expected, servers commit commands and create a new checkpoint that reflects the current state of servers. A checkpoint $C_{i}$ of server $i$ is a triple $(S,P,W)$, where $S$ is the current state of the state machine, $P$ is a sequence of sufficiently old (but not yet committed) commands, and $W$ is the $T$-window index when the checkpoint was taken. The parameters $S$ and $P$ represent a valid state of server $i$ at some point, while $W$ provides a priority ordering (where newer checkpoints take precedence). We also maintain a reset variable $R_{i}\in\{\texttt{reset},\texttt{no-reset},\bot\}$ for each server $i$, indicating whether $i$ intends to revert to a previous checkpoint. If $R_{i}=\texttt{no-reset}$, server $i$ expects normal operation and a benign adversary. If $R_{i}=\bot$, server $i$ believes a surge is currently in progress. If $R_{i}=\texttt{reset}$, server $i$ considers the surge over and that a rollback to a prior checkpoint is needed.
@@
The complete recovery protocol can be viewed as two separate protocols: the reset protocol and the checkpoint protocol that run alongside the extended median rule, and incorporates the functionality of the compact median rule with certificates. The reset protocol is governed by each server’s reset state, $R_{i}$. At the end of each $T$-window, if server $i$ has an empty log $L_{i}$ (from the extended median protocol), it concludes that it was blocked and sets $R_{i}=\texttt{reset}$. Otherwise, $R_{i}=\texttt{no-reset}$. During a $T$-window, in every round, each server $i$ queries the reset states $R_{j}$ of $k$ randomly chosen servers. If at least $\ell$ responses are received and any of them is no-reset, the server sets $R_{i}=\texttt{no-reset}$; if all responses are reset, it sets $R_{i}=\texttt{reset}$. If fewer than $\ell$ responses are received, $R_{i}$ becomes $\bot$, and a server in the $\bot$ state does not answer queries about its own state $R_{i}$. This design ensures that if the adversary surges during a $T$-window, the system falls into a “death spiral” where all servers eventually switch to $R_{i}=\bot$. Otherwise, the servers converge on either reset or no-reset. Finally, if a server ends a $T$-window with $R_{i}=\texttt{reset}$, it rolls back to a previously agreed-upon checkpoint.
@@
The checkpoint protocol revolves around the triple $C_{i}$. Servers exchange checkpoints by issuing $k$ random queries, analogous to how they exchange logs $\textit{L}_{i}$ or states $R_{i}$. The key distinction is that checkpoints are prioritized by age, given by the third parameter $W$. If the newest received checkpoint $(S^{\prime},P^{\prime},W^{\prime})$ has a higher number $W^{\prime}$ than its own checkpoint, it immediately adopts it and updates its state $S_{i}:=S^{\prime}$ and its log $L_{i}:=P^{\prime}$, since the server is outdated. This procedure spreads the newest checkpoints (those with the largest $W$) within one $T$-window, assuming $T=O(\log n)$ is sufficiently large.
@@
For conciseness, we assume in this section that the functionality of the compact median protocol with certificates is subsumed in the state machine, in particular, the required data structure centered around client sequence numbers (Section 4) and the Merkle hash forests (Section 5). The techniques shown in sections 4 and 5 can be adapted to work in the recovery protocol. Consequently, we will assume here that client commands are unique.
@@
    • $W$: number of $T$-window (initially 0)
@@ 7 Conclusion
We introduced a lightweight solution for state machine replication under an adversary capable of blocking any number servers. As shown in [DoerrGMSS11], the median rule can tolerate $\smash{O}(\sqrt{n})$ adversarial servers, suggesting that our approach can be extended to handle not just blocked but also $\smash{O}(\sqrt{n})$ malicious servers. Furthermore, we anticipate that it can accommodate server churn by allowing the adversary to have a $\beta$-fraction of blocked, joining or leaving servers in each round, thus removing the requirement of a static server set.
Our $1/10$ bound for blocked servers to ensure liveness is not fundamental. By tuning parameters $k,\ell$, the protocol tolerates any 1-late $(1-\varepsilon)$-blocking adversary for constant $\varepsilon>0$. However, our choice of parameters provides some measure of partition tolerance. With our chosen parameters, safety even holds under any partition as all but at most one component will “spiral to death” (see Lemma 2.3). This provides a trade-off in the sense of the CAP theorem between safety, liveness, and partition tolerance [GilbertL02].
Moreover, we believe our protocols can be extended so that commitment errors are not persistent. First of all, we ensure that commitment errors are very unlikely and, therefore, rare in the first place. Second, the proof of Theorem 3 implies that the extended median rule achieves agreement on the position of a command $x$ within $O(\log n)$ rounds, w.h.p., irrespective of how the logs look like initially, i.e., commitment errors only have limited effects on other commands. Third, our recovery protocol only takes another checkpoint of a shared state every $T=\Theta(\log n)$ rounds so that there is sufficient time to arrive at a consensus checkpoint that resolves disagreement on the shared state caused by a commitment error, by making use of the median rule. This pushes the probability of a persistent error to a point where it can be truly ignored.
@@ Appendix A Probabilistic Concepts
In this section, we are going to cover the probabilistic concepts and inequalities that we use frequently throughout this paper. The following two inequalities will turn out to be useful (see [McDiarmid1998]).
@@
Finally, we state a technical lemma from from [DoerrGMSS11] (Claim 3.4), needed to show that in a $(k,\ell)$-gossip-rule, if only a single server initially has a broadcast value, then within $O(\log n)$ rounds, it will either spread to at least $c\log n$ servers or no server holds that value anymore.